OpenStudy (anonymous):

In a laboratory, two bottles contain mixture of acid and water in the ratio 2 : 5 in the first bottle and 7 : 3 in the second. The ratio in which the contents of these two bottles be mixed such that the new mixture has acid and water in the ratio 2 : 3 is (A) 4 : 15 (B)9 : 8 (C) 21 : 8 (D)1 : 2

OpenStudy (anonymous):

Cant seem to get it

OpenStudy (shubhamsrg):

in 1st bottle, every unit contains 2/7 acid and 5/7 water in 2nd bottle, every unit contains 7/10 acid and 3/10 water lets take p units from 1st bottle and q units from 2nd bottle and mix together, thus we have on mixing 2p/7 + 7q/10 acid and 5p/7 + 3q/10 water we need to find p/q where we are given that (2p/7 + 7q/10) / (5p/7 + 3q/10) = 2/3 let p/q = x so we are left with on simplifying, (20 x + 49)/(50x + 21) = 2/3 60x + 147 = 100x + 42 or 40x = 105 or x = 21/8 <--------ans..

OpenStudy (anonymous):

Thank you :) @shubhamsrg

OpenStudy (shubhamsrg):

glad to have helped :)

OpenStudy (anonymous):

How did you simplify down to find p/q?

OpenStudy (shubhamsrg):

took q common from both numerator and denominator , or you can also see that as substituting p = xq

OpenStudy (anonymous):

ofcourse

OpenStudy (shubhamsrg):

you asked something else? @Skaematik